A particle of mass 5kg moves along x-axis under the influence of two forces;force of attraction to the origin which is numerically equal to 40 times the instantaneous distance from the origin and damping force is proportional to instantaneous speed; when the speed is 10m/s the damping is 200. Assuming that the particle from rest at a distance of 20cm from the origin then

a)Write the differential equations

b)Find the position at any given time t

c) Determine the amplitude, frequency and period of the damping oscillation.




1
Expert's answer
2021-02-24T06:59:35-0500

Solution:

(a):

Given, when the speed is 10m/s the damping is 200 N.

Constant of proportionality "\dfrac{200}{10}=20" N-s/m

So, making an equation using given forces as follows:

"5a=-40x-20v" , where a is acceleration, x is displacement and v is velocity.

"\Rightarrow a+4v+8x=0
\\ \Rightarrow \dfrac{d^2x}{dt^2}+4\dfrac{dx}{dt}+8x=0"

Required differential equation is "\dfrac{d^2x}{dt^2}+4\dfrac{dx}{dt}+8x=0"

(b):

Now, solving this differential equation to get its position.

Its auxiliary equation is "D^2+4D+8=0"

"\Rightarrow D=\dfrac{-4\pm\sqrt{4^2-4(8)}}{2}=\dfrac{-4\pm\sqrt{-32}}{2}=\dfrac{-4\pm4i\sqrt{2}}{2}"

"\Rightarrow D=-2\pm2\sqrt{2}i"

Thus, its general solution is: "x(t)=e^{-2t}(c_1\cos 2\sqrt2t+c_2\sin 2\sqrt2t)"

Or "x(t)=Re^{-2t}\cos (2\sqrt2t-\theta)" ...(i)

Now, put "t=0,x=20\ cm=0.2\ m"

"0.2=Re^{0}\cos (0-\theta)
\\ \Rightarrow R\cos \theta=0.2" ...(ii)

Differentiating (i) w.r.t t, we get,

"x'(t)=(-2)Re^{-2t}\cos (2\sqrt2t-\theta)-Re^{-2t}\sin (2\sqrt2t-\theta)(2\sqrt2)"

Also, "\frac{dx}{dt}=0,\ when\ t=0" . Putting these values, we get,

"0=(-2)Re^{0}\cos (0-\theta)-Re^{0}\sin (0-\theta)(2\sqrt2)
\\ 0=(-2)R\cos (\theta)+R\sin (\theta)(2\sqrt2)"

"0=(-2)0.2+\dfrac{0.2}{\cos\theta}\sin (\theta)(2\sqrt2)" [Using (ii)]

"1=\tan\theta(\sqrt2)"

"\dfrac1{\sqrt2}=\tan\theta"

"\Rightarrow \theta=35.26\degree"

Put this in (ii)

"R\cos 35.26\degree=0.2"

"R(0.8165)=0.2"

"R=0.2449\approx0.25"

Hence, the position at any time t is: "x(t)=0.25e^{-2t}\cos (2\sqrt2t-35.26\degree)"

(c):

Amplitude"=0.25e^{-2t}\ m"

Period"=\dfrac{2\pi}{2\sqrt2}=\dfrac{\pi}{\sqrt2} \ s"

Frequency"=\dfrac{\sqrt2}{\pi} \ Hz"